Find the probability of selecting both red cards when two cards are drawn (without replacement) from a standard deck of cards.: Find the probability of selecting both red cards when two cards are drawn (without replacement) from a standard deck of cards.
Answer by checkley77(3397)

26/52*25/51=650/2,652=.245=24.5% probability of picking 2 red cards.